A sample of drinking water has a nitrate concentration 2.3 ppm (a level that is safe to drink). Calculate the mass, in grams, of nitrate ions in a 250 mL glass of this water.


1
Expert's answer
2021-03-22T06:38:37-0400

According to the definition of ppm (parts per million), 1 ppm = 0.001 g / 1000 mL = 1 mg/L.

From here, 250 mL of water contains:

m(nitrate) = 1 mg/L × 0.25 L × 2.3 ppm = 0.575 mg = 5.75 × 10-4 g


Answer: 5.75 × 10-4 g